OpenStudy (anonymous):

Find the extrema and points of inflection for f(x) = 3x / (x+8)^2

OpenStudy (anonymous):

\[3x / (x+8)^{2}\]

OpenStudy (bahrom7893):

so to do that yo will need to take the first derivative!

OpenStudy (bahrom7893):

I never remember the quotient rule off the top of my head so I will improvise.. haha

OpenStudy (anonymous):

dtop(bottom)-d(bottom)top=0 rtemeber you dont need to do bottom squared and set that equal to 0 because at 0 its undefined ;)

OpenStudy (bahrom7893):

f(x) = 3x * (x+8)^(-2) and just use the chain rule

OpenStudy (bahrom7893):

f '(x) = 3x * (-2)(x+8)^(-3) + (x+8)^(-2) * 3

OpenStudy (anonymous):

no this would be product rule not the chain

OpenStudy (bahrom7893):

woops sorry meant product and chain rules..

OpenStudy (anonymous):

Don't bother, bahrom7893, I've got it here (points to head). The problem is that with all my homework problems, I've been getting the wrong answer somehow. This was the same stuff I was nailing on my test last Friday, but it's coming out WRONG on homework? Quotient Rule: f(x) g'(x) - f'(x)g(x) / g(x)^2

OpenStudy (bahrom7893):

ohh do u just want the answer?

OpenStudy (anonymous):

if it's OK, I'd prefer that you guys watch the steps I take and let me know if I mess up.

OpenStudy (anonymous):

That is if you have some free time...

OpenStudy (anonymous):

shoot

OpenStudy (bahrom7893):

then ill continue: f '(x) = 3x * (-2)(x+8)^(-3) + (x+8)^(-2) * 3 f'(x) = [-6x/(x+8)^3] + [3/(x+8)^2]

OpenStudy (bahrom7893):

okay from now on typing is going to be a pain in the asymptote so let's just go on twiddla.. http://www.twiddla.com/575500

OpenStudy (anonymous):

It would take too long to write it out. For the first derivative, i got -3(x+16) / 2(x+8)^3

OpenStudy (anonymous):

bahrom7893 is getting -3(x+8)...I'm about to give up. LOL

OpenStudy (anonymous):

3x / (x+8)^2 f'=3(x+8)^2-(2(x+8)(3x)/(x+8)^4=0 (x^2+16x+64)-2x^2-16x=0 -x^2+64=0 (x-8)x(+8)=0 x=+-8

OpenStudy (anonymous):

So I shouldn't convert it to (x+8)^-2 for f(x) and use the product rule? Stick with the layout given by the problem (requiring quotient rule)?

OpenStudy (anonymous):

then take the f'' which looks nasty and i dont want to do it but its the same as the first part i did those numbers extrema and the f'' are either positive or negative and those are the concave up or down

OpenStudy (anonymous):

NOOOOOOOOOO yes use quoetient rule

OpenStudy (anonymous):

but at -8 its undefined so +8 is the max or min depending on f''

OpenStudy (anonymous):

for f'' you have to test numbers around the one that =0 remeber to see if it goes -,+,- so its concave down

OpenStudy (anonymous):

to do a simple one f(x)=x^3 f'=3x^2=0 x=0 so thats a min or max f''=6x=0 x also is 0 test-100,100 and at f''(-100) you get <0 f(100)=>0 so at 0 you get a min which is concave up which is exactly what you see in the graph

OpenStudy (bahrom7893):

seems i've been wasting time on twiddla... explaining all that to nobody..